In a survey, repondents were asked, How many years have you worked in the same job?. The relative frequency histogram is shown below. Relative Frequency 0.5 0.4 0.3 0.2 0.1 0-5 years 5- 10 years 10-15 years 15 -20 years (a) What percent of the population answered 10 or fewer years? % (a) If the population size was 200, how many people answered 15-20 years?
Answer #1

Answer : the given chart gives the relative frequency data.

(a) population that answered 10 or fewer years = population that answered between 0-10 years = 0.2+0.4 = 0.6

?age = 0.6*100 = 60%

(b) relative frequency of persons answering 15-20 years = 0.1

If population size is 200 than it would be = 200 * 0.1 = 20.

Please up vote.